2) Elasticity

Return to the demand curve from question (1), that is, P = -0.4QD + 120 a) What is the elasticity of demand going from P = 100 to P = 35? b) What is the elasticity of demand going from P = 35 to P = 100? c) Explain the discrepancy in the amounts

d) What is one alternative measurement method that addresses the discrepancy? Calculate the elasticity using this method.

e) Is demand elastic or inelastic? If the seller wanted to maximize revenue based just on the demand curve, should he/she raise or lower price? Why?
Answer #1

a) At P=100, Q =20/0.4=50, At P=35 Q=85/0.4=212.5 ,Price elasticity of demand using percentage change method=%change in demand/% change in price=((212.5-50)/50)/((35-100)/100)=-5

b) when price moved from 35 to 100 elasticity=((50-212.5)/212.5)/((100-35)/35)=-0.411

c) Discrepency is caused because of relative percentages in increase /decrease

d) The problem can be avoided using mid point method basis which elasticity is calculated for changes based average of the price/quantity. Elasticity here will be same for both the cases i.e ((212.5-50)/((212.5+50)/2))/((35-100)/((35+100)/2)=-1.29

e) Demand is elastic since |e|>1 hence if person wants to increase revenue he/she should lower the prices As the demand changes at faster rate for a unit change in price hence the quantity demanded hence demand will have more weight than price for revenue